# Learning Flatness-Preserving Residuals for Pure-Feedback Systems

### v2

@@ Abstract
We study residual dynamics learning for differentially flat systems, where a nominal model is augmented with a learned correction term from data. A key challenge is that generic residual parameterizations may destroy flatness, limiting the applicability of flatness-based planning and control methods. To address this, we propose a framework for learning flatness-preserving residual dynamics in systems whose nominal model admits a pure-feedback form. We show that residuals with a lower-triangular structure preserve both the flatness of the system and the original flat outputs. Moreover, we provide a constructive procedure to recover the flatness diffeomorphism of the augmented system from that of the nominal model. We then introduce a learning algorithm that fits such residuals from trajectory data using smooth function approximators. Our approach is validated in simulation on a 2D quadrotor subject to unmodeled aerodynamic effects. We demonstrate that the resulting learned flat model enables tracking performance comparable to nonlinear model predictive control (5×5$\times$5 × lower tracking error than the nominal flat model) while also achieving over a 20×20$\times$20 × speedup in computation.
1 Introduction
@@
Motivated by this shortcoming, we propose augmenting a differentially flat nominal model by learning to approximate the unmodeled (i.e., residual) dynamics from trajectory data. However, existing residual learning algorithms (Shi et al., 2019; Carron et al., 2019; Torrente et al., 2021; Chee et al., 2022) typically use generic functional forms for the residual, which can lead to augmented dynamics that no longer enjoy the flatness property. Furthermore, even if flatness is preserved, it is often unclear how to construct the flatness diffeomorphism for the resulting system, rendering flatness-based algorithms inapplicable. To overcome this, we restrict our attention to a class of differentially flat nominal dynamics that admit a pure-feedback form and derive a sufficient condition under which the residual preserves flatness. We show that when the residual is lower-triangular—a property that preserves the pure-feedback structure of the nominal dynamics—the flatness diffeomorphisms of the augmented system can be constructed using those of the nominal model. We validate our approach on planar quadrotor dynamics, achieving tracking performance comparable to a nonlinear model predictive controller (NMPC) while requiring an order of magnitude less computation time. To the best of our knowledge, this is the first work to address learning flatness-preserving residual dynamics for differentially flat systems.
1.1 Related Work
1.1.1 Learning Residual Dynamics
@@
1.1.2 Learning to Linearize Nonlinear Systems
Learning-based techniques have been applied to differentially flat systems to derive linearizing coordinate transformations from data. Representative examples include seeking to identify flatness diffeomorphisms from input-output data (Sferrazza et al., 2016; Ma et al., 2022), and directly learning a feedback-linearizing control law (Umlauft et al., 2017; Westenbroek et al., 2019; Greeff and Schoellig, 2020). For pure-feedback systems, adaptive control methods have also been proposed to deal with parametric uncertainty in the dynamics, including model reference adaptive control (Nam and Araposthathis, 1988) and neural control methods that leverage backstepping (Ge and Wang, 2002, 2004). In contrast to these approaches, which are complementary to our work, we focus on learning additive residual dynamics from data while ensuring that the resulting augmented model preserves differential flatness.
1.2 Contributions
@@
  • Learning algorithm: We develop a residual learning framework that enforces this structure, enabling flatness-preserving residual models to be learned directly from trajectory data.
  • Empirical validation: We demonstrate in simulation that our method matches the tracking performance of NMPC while achieving an order-of-magnitude improvement in computational efficiency.
2 Preliminaries and Problem Formulation
@@
Assumption 1 implicitly assumes that the $m$ sub-states $\mathbf{x}_{i,1},\ldots,\mathbf{x}_{i,m}$ have the same relative degree $r$, and hence all $\mathbf{x}_{i}$ share the same dimension with $\mathbf{u}$. We choose this formulation for ease of notation and exposition, but our analysis can be adapted to the more general case where the sub-states have different relative degrees.
@@
3 Differentially Flat Pure-Feedback Systems
@@
Many well-known flat systems (e.g., unicycles and fully-actuated manipulators) can be placed into a form that complies with Assumptions 1 and 2 under suitable dynamic extensions. [^2] The need for dynamic extensions can be avoided if we consider pure-feedback systems whose sub-states do not have uniform relative degrees (as described in Remark 1). Furthermore, while Theorem 1 characterizes local differential flatness around an operating point, many well-known flat systems are differentially flat over a large subset of the state and control spaces.
@@
$\dot{\mathbf{x}}_{1}=\bar{f}_{1}(\mathbf{x}_{2})=\mathbf{x}_{2},$ $\dot{\mathbf{x}}_{2}=\bar{f}_{2}(\mathbf{x}_{3})$ $=\frac{1}{\mathfrak{m}}\begin{bmatrix}-F\sin\theta\\ F\cos\theta-\mathfrak{m}g\end{bmatrix},$
$\dot{\mathbf{x}}_{3}=\bar{f}_{3}(\mathbf{x}_{4})=\mathbf{x}_{4},$ $\dot{\mathbf{x}}_{4}=\bar{f}_{4}(\mathbf{u})$ $=\begin{bmatrix}1&0\\ 0&I^{-1}\end{bmatrix}\mathbf{u}.\qquad$ (7)
@@
4 Flatness-Preserving Perturbations
@@
While this structure may seem overly restrictive, we highlight that particular instances satisfying these assumptions have already been identified in the literature, and we highlight one such example in the context of unmodeled aerodynamics for aerial vehicles at the end of this section.
@@
$=\begin{bmatrix}\mathfrak{m}\sqrt{(\ddot{\mathbf{y}}_{1}+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{1})^{2}+(\ddot{\mathbf{y}}_{2}+g+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{2})^{2}}\\ \arctan 2(-\ddot{\mathbf{y}}_{1}+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{1},\ddot{\mathbf{y}}_{2}+g+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{2})\end{bmatrix}.$
@@
5 Learning Flatness-Preserving Residuals
The results in Section 4 identify a rich class of residual dynamics (Def. 2) that can be added to a flat nominal system in pure-feedback form so as to define augmented systems that are also differentially flat (Thm. 2), and for which the corresponding flatness diffeomorphisms can be easily constructed given those of the nominal system (Thm. 3). We leverage these insights to propose an algorithm for learning flatness-preserving residuals from data.
@@ 5 Learning Flatness-Preserving Residuals
where $\Theta:=(\theta_{1},\ldots,\theta_{r})$ are parameters of suitable dimensions and the $\hat{\Delta}_{i}$ are continuously differentiable in each $\mathbf{x}_{j}$. By Theorem 2, the augmented system $\bar{f}+\hat{\Delta}_{\Theta}$ is differentially flat, and its flatness diffeomorphisms can be constructed according to equation (10). A simple and practically appealing choice for the $\hat{\Delta}_{i}$ is to parameterize them as neural networks with continuously differentiable activation functions.
@@
$L(\Theta)=\sum_{i=1}^{N}\sum_{k=1}^{T-1}\left\lVert\Delta(\boldsymbol{\tau}^{i}_{k};\Theta)-\hat{\Delta}(\boldsymbol{\tau}^{i}_{k};\Theta)\right\rVert_{2}^{2},$ (11)
@@
$\Delta(\boldsymbol{\tau}^{i}_{k}):=\frac{\mathbf{x}^{i}(t_{k+1})-\mathbf{x}^{i}(t)}{t_{k+1}-t_{k}}-\bar{f}(\mathbf{x}^{i}(t_{k}),\mathbf{u}^{i}(t_{k})).$
@@
6 Numerical Experiments
We validate our algorithm via simulation experiments on the 2D quadrotor (7) under unmodeled air resistance. First, we test whether augmenting the nominal model with learned residual dynamics can generate more dynamically consistent trajectories in an open-loop rollout (Section 6.2). Then, we design a flatness-based tracking controller using the learned model and compare its tracking performance and computation time against an NMPC baseline (Section 6.3). We show that the flatness-based controller achieves tracking performance similar to that of the NMPC, while requiring an order of magnitude less time to compute the control action. This computational efficiency will allow it to be run at a higher frequency and makes it more suitable for computationally constrained hardware. Code to reproduce our experiments is available at https://github.com/FJYang96/flat-residual/.
6.1 Setup
@@
6.2 Open-Loop Trajectory Evaluation
We first test the augmented model’s ability to generate dynamically consistent state and input trajectories from trajectories of the flat outputs. For each reference trajectory ($\mathbf{p}_{circ}$ or $\mathbf{p}_{lem}$), we generate open-loop control inputs in two ways: (1) using the flatness diffeomorphism, $(\bar{\Phi},\bar{\Psi})$, of the nominal 2D quadrotor dynamics model $\bar{f}$ (which ignores drag), and (2) using the augmented diffeomorphism, $(\hat{\Phi},\hat{\Psi})$, of the augmented dynamics model $\bar{f}+\hat{\Delta}_{\Theta}$ that accounts for drag. We then roll out the open loop control actions under the true dynamics $f$. Recall that for our system (7), the flat outputs are the position (i.e., ${\mathbf{y}=\mathbf{x}_{1}=\mathbf{p}}$). For the nominal diffeomorphism, we compute the trajectory $\bar{\mathbf{x}}(t)$ that satisfies
@@ 6.2 Open-Loop Trajectory Evaluation
for the augmented system. In both cases, we integrate these ODE’s numerically using the Runge-Kutta method with a step size of $10^{-2}$ seconds. We report the position error, computed for a trajectory $\mathbf{x}(t)$ as
$\mathbf{e}=\frac{1}{k}\sum_{i=1}^{T}\left\lVert\mathbf{x}_{1}(t_{k})-\mathbf{p}(t_{k})\right\rVert$
in Table 1, and show representative trajectories in Figure 2. We note that, on average, incorporating the learned residual model reduces position error by nearly an order of magnitude. This suggests that the learned flat model is able to capture the unmodeled effects and adjust the flatness diffeomorphisms accordingly, showing its potential to be used for accurate trajectory planning. We also repeated the same steps for a flatness diffeomorphism $(\Phi^{*},\Psi^{*})$ constructed with the ground-truth residual. Consistent with our theoretical results, the corresponding simulated open-loop trajectory has a position error less than $10^{-5}$, providing further evidence that the procedure in (10) correctly constructs the diffeomorphism and thus corroborating our claims in Theorem 3.
@@
Table 1: Mean Open-Loop Position Error (m) $\pm$ Standard Deviation over $30$ training runs.
6.3 Closed-Loop Tracking Control
@@
[…]
@@
The controller then applies the control action $\mathbf{u}(0)$ until the next sampling step. In practice, we solve this problem using the direct method. We consider a prediction horizon of ${H=1}$ second and discretize the dynamics into $100$ time steps using RK4 integration. We use the cost matrices $Q=Q_{T}=I$ and $R=0.001I$. We warmstart the optimization problem from the previous solution and solve it in Python using Casadi (Andersson et al., 2019) until convergence. We also assume that the controller observes the full state at each time step.
@@
Figure 3: Closed-Loop Tracking Control Results: Figs. 3(a) and 3(b) show representative trajectories tracking the circular and lemniscate references, and Fig. 3(c) shows mean and standard deviation of compute time per control update for both methods. The flatness-based controller with learned residual dynamics achieves similar performance to nonlinear MPC while requiring 20×20$\times$20 × less time to compute control actions.
@@
Table 2: Mean Closed-loop Tracking Error (m) $\pm$ Standard Deviation over $30$ training runs.
For both methods, the control actions sampled with zero-order hold are updated at a rate of $100Hz$. Like the open-loop case, we report the position error in Table 2 and show representative trajectories in Figure 3. We observe that our learned flatness-based controller is able to perform comparably to NMPC, with both showing more than 5×5$\times$5 × less position tracking error than the nominal flatness-based controller. Further, despite the need to compute Jacobians of the neural networks, the simple structure of the flatness-based controller allows it to be much more computationally efficient. We show the results on computation time in Figure 3(c). Specifically, the flatness-based controller averages $3.3\pm 0.3ms$ of wall-clock time per control action update. In contrast, even with warm-starting, NMPC takes $85.2\pm 1.74ms$. This suggests the potential to run the learned flatness-based controller at a significantly higher frequency or on much more computationally constrained hardware. Although our implementation for both methods has room for optimization, the observed computational savings are in line with those observed in prior work (Sun et al., 2022), wherein the authors show that flatness-based controllers are able to run 2 orders of magnitude faster than NMPC in real-world hardware experiments, with similar tracking performance.
7 Conclusion

### v3

@@ Abstract
We study residual dynamics learning for differentially flat systems, where a nominal model is augmented with a learned correction term from data. A key challenge is that generic residual parameterizations may destroy flatness, limiting the applicability of flatness-based planning and control methods. To address this, we propose a framework for learning flatness-preserving residual dynamics in systems whose nominal model admits a pure-feedback form. We show that residuals with a lower-triangular structure preserve both the flatness of the system and the original flat outputs. Moreover, we provide a constructive procedure to recover the flatness diffeomorphism of the augmented system from that of the nominal model. Building on these insights, we introduce a parameterization of flatness-preserving residuals using smooth function approximators, making them learnable from trajectory data with conventional algorithms. Our approach is validated in simulation on a 2D quadrotor subject to unmodeled aerodynamic effects. We demonstrate that the resulting learned flat model achieves a tracking error $5\times$ lower than the nominal flat model, while being $20\times$ faster over a structure-agnostic alternative.
I Introduction
@@
Motivated by this shortcoming, we propose augmenting a differentially flat nominal model by learning to approximate the unmodeled (i.e., residual) dynamics from trajectory data. However, existing residual learning algorithms (Shi et al., 2019; Carron et al., 2019; Torrente et al., 2021; Chee et al., 2022) typically use generic functional forms for the residual, which can lead to augmented dynamics that no longer enjoy the flatness property. Furthermore, even if flatness is preserved, it is often unclear how to construct the flatness diffeomorphism for the resulting system, rendering flatness-based algorithms inapplicable. To overcome this, we restrict our attention to a class of differentially flat nominal dynamics that admit a pure-feedback form and derive a sufficient condition under which the residual preserves flatness. We show that when the residual is lower-triangular—a property that preserves the pure-feedback structure of the nominal dynamics—the flatness diffeomorphisms of the augmented system can be constructed using those of the nominal model. We validate our approach on planar quadrotor dynamics, achieving significantly lower tracking error than without learning while requiring an order of magnitude less computation time than a structure-agnostic approach that incorporated the learned residual via nonlinear model predictive control (NMPC). To the best of our knowledge, this is the first work to address learning flatness-preserving residual dynamics for differentially flat systems.
I-A Related Work
I-A1 Learning Residual Dynamics
@@
I-A2 Learning to Linearize Nonlinear Systems
Learning-based techniques have been applied to differentially flat systems to derive linearizing coordinate transformations from data. Representative examples include seeking to identify flatness diffeomorphisms from input-output data (Sferrazza et al., 2016; Ma et al., 2022), and directly learning a feedback-linearizing control law (Umlauft et al., 2017; Westenbroek et al., 2019; Greeff and Schoellig, 2020). For pure-feedback systems, adaptive control methods have also been proposed to deal with parametric uncertainty in the dynamics, including model reference adaptive control (Nam and Araposthathis, 1988) and neural control methods that leverage backstepping (Ge and Wang, 2002, 2004). In a related vein, model-free control techniques exist for flat systems to estimate and reject more general disturbances online (Fliess and Join, 2013; Join et al., 2024). In contrast to these approaches, which are complementary to our work, we focus on learning residual dynamics from data while ensuring that the resulting augmented model preserves flatness.
I-B Contributions
@@
  • Learning algorithm: We develop a parameterization of the residual dynamics that enforces this structure, enabling flatness-preserving residual models to be learned directly from trajectory data.
  • Empirical validation: We demonstrate in simulation that our method significantly reduces tracking error relative to the nominal controller while being an order-of-magnitude more computationally efficient than an NMPC-based structure-agnostic approach.
II Preliminaries and Problem Formulation
@@
Assumption 1 implicitly assumes that the $m$ sub-states $\mathbf{x}_{i,1},\ldots,\mathbf{x}_{i,m}$ have the same relative degree $r$, and hence all $\mathbf{x}_{i}$ share the same dimension with $\mathbf{u}$. We choose this formulation for ease of notation and exposition, but our analysis can be adapted to the case where the sub-states have different relative degrees via dynamic extension .
@@
III Differentially Flat Pure-Feedback Systems
@@
Many well-known flat systems (e.g., unicycles and fully-actuated manipulatorsRathinam and Murray (1998)) can be placed into a form that complies with Assumptions 1 and 2 under suitable dynamic extensions. Furthermore, while Theorem 1 characterizes local differential flatness around an operating point, many well-known flat systems are differentially flat over a large subset of the state and control spaces.
@@
$\dot{\mathbf{x}}_{1}=\bar{f}_{1}(\mathbf{x}_{2})=\mathbf{x}_{2},\ \$$\dot{\mathbf{x}}_{2}=\bar{f}_{2}(\mathbf{x}_{3})$ $=\frac{1}{\mathfrak{m}}\begin{bmatrix}-F\sin\theta\\ F\cos\theta-\mathfrak{m}g\end{bmatrix}$,
$\dot{\mathbf{x}}_{3}=\bar{f}_{3}(\mathbf{x}_{4})=\mathbf{x}_{4},\ \$$\dot{\mathbf{x}}_{4}=\bar{f}_{4}(\mathbf{u})$ $=\begin{bmatrix}1&0\\ 0&I^{-1}\end{bmatrix}\mathbf{u}.\qquad$ (7)
@@
IV Flatness-Preserving Perturbations
@@
While this structure may seem overly restrictive, particular instances satisfying these assumptions have already been identified in the literature. We highlight one such example in the context of unmodeled aerodynamics for aerial vehicles at the end of this section.
@@
$=\begin{bmatrix}\mathfrak{m}\sqrt{(\ddot{\mathbf{y}}_{1}+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{1})^{2}+(\ddot{\mathbf{y}}_{2}+g+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{2})^{2}}\\ {\tan^{-1}\left(\left(\ddot{\mathbf{y}}_{2}+g+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{2}\right)/\left(-\ddot{\mathbf{y}}_{1}+\frac{C_{r}}{\mathfrak{m}}\dot{\mathbf{y}}_{1}\right)\right)}\end{bmatrix}$.
@@
V Learning Flatness-Preserving Residuals
The results in Section IV identify a rich class of residual dynamics (Def. 2) that can be added to a flat nominal system in pure-feedback form so as to define augmented systems that are also differentially flat (Thm. 2), and for which the corresponding flatness diffeomorphisms can be easily constructed given those of the nominal system (Thm. 3). We leverage these insights to propose an algorithm for learning flatness-preserving residuals from data.
@@ V Learning Flatness-Preserving Residuals
where $\Theta:=(\theta_{1},\ldots,\theta_{r})$ are parameters of suitable dimensions and the $\hat{\Delta}_{i}$ are continuously differentiable in each $\mathbf{x}_{j}$. By Theorem 2, the augmented system $\bar{f}+\hat{\Delta}_{\Theta}$ is differentially flat, and its flatness diffeomorphisms can be constructed according to equation (10). A practically appealing and standard choice is to parameterize the $\hat{\Delta}_{i}$ as neural networks with continuously differentiable activation functions.
@@
$L(\Theta)=\sum_{i=1}^{N}\sum_{k=1}^{T-1}\left\lVert\Delta(\boldsymbol{\tau}^{i}_{k})-\hat{\Delta}(\boldsymbol{\tau}^{i}_{k};\Theta)\right\rVert_{2}^{2},$ (11)
@@
$\Delta(\boldsymbol{\tau}^{i}_{k}):=\frac{\mathbf{x}^{i}(t_{k+1})-\mathbf{x}^{i}(t_{k})}{t_{k+1}-t_{k}}-\bar{f}(\mathbf{x}^{i}(t_{k}),\mathbf{u}^{i}(t_{k})).$
@@
VI Numerical Experiments
We validate our algorithm via simulation experiments on the 2D quadrotor (7) under unmodeled air resistance. First, we test whether augmenting the nominal model with learned residual dynamics can generate more dynamically consistent trajectories in an open-loop rollout (Section VI-B). Then, we design a flatness-based tracking controller using the learned model and compare its tracking performance and computation time against an NMPC baseline (Section VI-C). We show that the flatness-based controller achieves tracking performance similar to that of the NMPC, while requiring an order of magnitude less time to compute the control action. This computational efficiency will allow it to be run at a higher frequency and makes it more suitable for computationally constrained hardware. Code to reproduce our experiments is available at https://github.com/FJYang96/flat-residual/.
VI-A Setup
@@
VI-B Open-Loop Trajectory Evaluation
[…]
@@ VI-B Open-Loop Trajectory Evaluation
for the augmented system. In both cases, we integrate these ODE’s numerically using the Runge-Kutta method with a step size $\tau=10^{-2}$ seconds. We report the position error, computed for a trajectory $\mathbf{x}(t)$ as
$\mathbf{e}=\frac{1}{\lceil T/\tau\rceil}\sum_{k=1}^{\lceil T/\tau\rceil}\left\lVert\mathbf{x}_{1}(k\tau)-\mathbf{p}(k\tau)\right\rVert$
in Table I, and show representative trajectories in Figure 2. We note that, on average, incorporating the learned residual model reduces position error by nearly an order of magnitude. This suggests that the learned flat model is able to capture the unmodeled effects and adjust the flatness diffeomorphisms accordingly, showing its potential to be used for accurate trajectory planning. We also repeated the same steps for a flatness diffeomorphism $(\Phi^{*},\Psi^{*})$ constructed with the ground-truth residual. Consistent with our theoretical results, the corresponding simulated open-loop trajectory has a position error less than $10^{-5}$, providing further evidence that the procedure in (10) correctly constructs the diffeomorphism and thus corroborating our claims in Theorem 3.
@@
TABLE I: Mean Open-Loop Position Error (m) $\pm$ Standard Deviation over $30$ training runs.
VI-C Closed-Loop Tracking Control
@@
$\underset{\mathbf{x},\mathbf{u}}{\text{minimize}}$ $\quad\int_{t=0}^{H}\Big((\mathbf{y}-\mathbf{y}_{r})^{\top}Q(\mathbf{y}-\mathbf{y}_{r})+\mathbf{u}^{\top}R\mathbf{u}\Big)dt$
@@
The controller then applies the control action $\mathbf{u}(0)$ until the next sampling step. In practice, we solve this problem using the direct method. We consider a prediction horizon of H=1{H=1} second and discretize the dynamics into $100$ time steps using RK4 integration. We use the cost matrices $Q=Q_{T}=I$ and $R=0.001I$. We warmstart the optimization problem from the previous solution and solve it in Python using Casadi (Andersson et al., 2019) until convergence. We also assume that the controller observes the full state at each time step.
@@
Figure 3: Closed-Loop Tracking Control Results: Figs. 3(a) and 3(b) show representative trajectories tracking the circular and lemniscate references, and Fig. 3(c) shows mean and standard deviation of compute time per control update for both methods. The flatness-based controller with learned residual dynamics achieves similar performance to nonlinear MPC while requiring $20\times$ less time to compute control actions.
@@
TABLE II: Mean Closed-loop Tracking Error (m) $\pm$ Standard Deviation over $30$ training runs.
For both methods, the control actions sampled with zero-order hold are updated at a rate of $100Hz$. Like the open-loop case, we report the position error in Table II and show representative trajectories in Figure 3. We observe that by incorporating the learned residual model, both the flatness-based controller and the NMPC showed more than $5\times$ less position tracking error than the nominal flatness-based controller. However, despite the need to compute Jacobians of the neural networks, the simple structure of the flatness-based controller allows it to be much more computationally efficient. We show the results on computation time in Figure 3(c). Specifically, the flatness-based controller averages $3.3\pm 0.3ms$ of wall-clock time per control action update. In contrast, even with warm-starting, NMPC takes $85.2\pm 1.74ms$. This suggests the potential to run the learned flatness-based controller at a significantly higher frequency or on much more computationally constrained hardware. Although our implementation for both methods has room for optimization, the observed computational savings are in line with those observed in prior work (Sun et al., 2022), wherein the authors show that flatness-based controllers are able to run 2 orders of magnitude faster than NMPC in real-world hardware experiments, with similar tracking performance.
VII Conclusion